# Robust shallow shadows

### v1

@@ Abstract
We present a robust shadow estimation protocol for wide classes of shallow measurement circuits that mitigates noise as long as the effective measurement map is locally unitarily invariant. This is in practice an excellent approximation, encompassing for instance the case of ideal single-qubit Clifford gates composing the first circuit layer of an otherwise arbitrary circuit architecture and even non-Markovian, gate-dependent noise in the rest of the circuit. We argue that for approximately local noise the measurement channel has an efficient matrix-product (tensor-train) representation, and show how to estimate this directly from experimental data using tensor-network tools, eliminating the need for analytical or numeric calculations. We illustrate the relevance of our method with both numerics and proof-of-principle experiments on an IBM Q device. Numerically, we show that, while unmitigated shallow shadows with noisy circuits become more biased as the depth increases, robust ones become more accurate for relevant parameter regimes. Experimentally, we observe major bias reductions in two simple fidelity estimation tasks using 5-qubit circuits with up to 2 layers of entangling gates using the mitigated variant, of close to an order of magnitude for $10^{4}$ measurement shots, e.g. Under the practical constraints of current and near-term noisy quantum devices, our method maximally realizes the potential of shadow estimation with global rotations.
@@ Introduction.
Randomized measurements are a central tool for the estimation of observables on repeatedly prepared quantum states [1, 2, 3]. Applications abound [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Ref. [16] derived guarantees and proved optimality of linear estimation from bases measurements generated by global or local Clifford rotations and coined the term “classical shadows”. Besides optimality, an important contribution of Ref. [16] is the efficiency of estimation from global Clifford rotations. For example, with the same classical shadow from constantly many state copies in the system size $n$, one can estimate many different pure state fidelities. In practice, however, harnessing this efficiency is typically not possible, since global Clifford rotations require the accurate implementation of “long” circuits of $\mathcal{O}(n^{2}/\log(n))$ many local gates [17]. This obstacle has motivated the study of “shallow shadows”, where shorter circuits are employed to “interpolate” between local and global rotations [18, 19, 20, 21, 22].
@@
While the two-design property of uniformly random Clifford rotations allows for straight-forward analytical inversion of the effective measurement map [23], a prerequisite to construct the classical shadow, the works on shallow circuits resort to numerical evaluation of the effective measurement map for depth larger than two [21]. However, for measurement circuits involving entangling gates, noise and (calibration) errors will typically cause sizeable biases in the derived estimators on current noisy intermediate-scale quantum (NISQ) hardware. For this reason, using shallow circuits for precise estimations in practice unavoidably requires a variant of classical shadows that mitigates circuit noise. That such a mitigation is possible for classical shadows with single-qubit or global random Clifford rotations if the noise is well-characterized was shown for simple depolarizing noise models [24], local gate-dependent noise models [25], and via randomized benchmarking experiments [26]. Going beyond these works, ideally one wishes a robust shadow scheme that replaces analytical or numerical calculations of the measurement map by a direct empirical estimation of it in the presence of noise. Such a tool was developed in Ref. [27], again for the cases of single-qubit or global random Clifford rotations under the assumptions of gate-independent Markovian after-gate noise. However, an extension to the most crucial, experimentally-relevant regime of shallow circuits, where no analytical calculation of the measurement frame is in general available, is missing.
@@
We explore this territory. We present a robust shadow estimation scheme for shallow circuits with arbitrary architectures of random local gates, under mild noise-type assumptions. We assume only that the average measurement channel is invariant under single-qubit Clifford rotations and accurate implementation of a single initial state. A simple but experimentally-crucial observation is that, for this to happen, it is sufficient for instance that the first circuit layer consists of ideal single-qubit random Cliffords, regardless of the subsequent circuit, which can even be subject to (time-stationary) gate-dependent, non-Markovian noise. Single-qubit Clifford invariance guarantees that the measurement map is Pauli diagonal. In turn, the circuit shallowness guarantees that the corresponding diagonal elements admit a tensor train (TT) description of low-rank if the noise is only weakly correlated. This allows us to develop efficient methods for directly estimating the measurement channel using standard tensor-network tools such as the modified alternating least squares ($\operatorname{MALS}$) [28].
@@
We numerically quantify how unmitigated shadow estimations (of both fidelities and Pauli observables) in the presence of noise become increasingly biased as the circuit depth increases. In contrast, we show that, for experimentally-relevant regimes, robust shallow shadows become more precise as the depth increases up to an optimal depth depending on the strength of the incoherence noise. For instance, for depth $D=2$, we observe average fidelity estimation errors over random target states almost one order of magnitude larger for the unmitigated estimations than for the robust ones already at $n=8$ qubits. Moreover, we apply our scheme to a recent experiment [5] where ideal, 4-qubit global Cliffords were approximated by depth-5 noisy circuits, in an unmitigated fidelity estimation. In numerical simulations, we observe a four-fold infidelity reduction with our method. Finally, we perform a proof-of-principle experiment on the ibm_nairobi device. We estimate the measurement frame on 5-qubit circuits of up to 2 layers of entangling gates, confirming significant deviations relative to the noiseless one. With this, we compare mitigated and unmitigated shadows in two simple fidelity estimation tasks. We observe major bias reductions using the robust variant, of close to an order of magnitude for $10^{4}$ measurements, e.g. Our framework offers a versatile and practical solution to classical shadows in the most relevant regime for near-term quantum devices, where randomized measurements are implemented by noisy shallow circuits.
Note that a related line of research studies the problem of mitigating the noise in the state-preparation using non-linear shadow estimation [29, 30, 9, 31] or existing error-mitigation techniques [32].
@@
We start by reviewing the classical shadow protocol on an ideal noiseless device. Consider a quantum device that can be repeatedly prepared in an unknown $n$-qubit quantum state $\varrho$. Furthermore, we will work with a quantum device that is capable of implementing circuits composed from a unitary gate set $\mathcal{G}$ and of performing a measurement in a fixed (computational) basis, specified by the effects =$\mathcal{B}=\{\Pi_{z}=\outerproduct{z}{z}\}_{z\in[d]}$, with $d=2^{n}$ and $[d]=\{0,1,\dots,d-1\}$.
@@
Every iteration of the protocol is described by the same positive operator-valued measure (POVM) taking values $\Pi_{z,g}=g^{\dagger}\,\Pi_{z}\,g$ with $g$ in the support of $\nu$ and $z\in[d]$. If the support of $\nu$ is sufficiently large, the POVM is informationally-complete (IC) and asymptotically any information about $\varrho$ can be reconstructed from the measurement statistic $\mathcal{S}$.
@@
where we have used Dirac notation with respect to the Hilbert-Schmidt inner product. Since the POVM is informationally complete (IC), $S$ is invertible and the elements of the canonical dual frame are given by $\bar{\Pi}_{z,g}=S^{-1}(\Pi_{z,g})$, fulfilling $\mathbb{E}_{g\sim\nu}{\big{[}}\sum_{z\in[d]}|\bar{\Pi}_{z,g})(\Pi_{z,g}|{\big{]}}=S^{-1}S=\mathbbm{1}_{d^{2}}$, with $\mathbbm{1}_{d^{2}}$ being the $d^{2}\times d^{2}$ identity operator. In the classical post-processing, we evaluate the scalar function $o(z,g)=(O|\bar{\Pi}_{z,g})$ on each entry of $\mathcal{S}$ and take the empirical mean $\hat{o}=\sum_{(z,g)\in\mathcal{S}}o(z,g)$. For the ideal device, we observe $(z,g)$ with probability density $p(z,g)=(\Pi_{z,g}|\varrho)\,\mathrm{d}\nu(g)$ and, thus, $\hat{o}$ is indeed an unbiased estimator of
@@
In practice, however, experimental imperfections in the implementation of the basis measurement $\mathcal{B}$ and the circuits in the support of $\nu$ induce a bias in the estimator. We denote by $\tilde{\Pi}_{z,g}$ the elements of the POVM that is actually implemented on the device and define the “one-sided” noisy frame operator as
@@
The same argument above now yields a non-vanishing estimator bias such that $\operatorname{bias}(O,\varrho)=|\mathbb{E}(\hat{o})-(O|\varrho)|=|(O|S^{-1}\tilde{S}-\operatorname{Id}_{d^{2}}|\varrho)|$.
The form of the bias already hints at a conceptionally simple solution to mitigate the noise. Assume we have an empirical estimate of the noisy frame operator $\tilde{S}$ and it is still invertible, which happens if $\{\tilde{\Pi}_{z,g}\}_{z,g}$ is still informationally complete. We can make use of this estimate in the construction of a scalar function $\tilde{o}(z,g)=(O|\tilde{S}^{-1}|\Pi_{z,g})$ that we evaluate instead of $o(z,g)$ in the classical post-processing. The resulting mean estimator $\hat{\tilde{o}}$ is again unbiased with $\mathbb{E}(\hat{\tilde{o}})=(O|\tilde{S}^{-1}\,\tilde{S}|\varrho)$.
@@
More concretely, recall that conjugation with the Clifford group (as with the unitary group) acts irreducibly on the space of traceless matrices, e.g. [2]. Correspondingly, tensor-products of single-qubit Clifford channels decompose into $d$ irreducible representations (irreps) without multiplicities labelled by $k\in\{0,1\}^{n}$. The projectors onto the irreps are $\Phi_{k}=\bigotimes_{l\in[n]}\Phi_{k_{l}}$ with $\Phi_{0}=d^{-2}\,|\mathbbm{1}_{2})(\mathbbm{1}_{2}|$ and $\Phi_{1}=\mathbbm{1}_{4}-\Phi_{0}$. Assuming single-qubit Clifford-invariance, the noisy frame operator can, thus, be written as
@@
Ref. [27] proposed the following calibration procedure to learn the coefficients $f_{k}$: Perform a randomized bases measurement on the input state |$\outerproduct{0}{0}$ |. The post-processing is analogous to that for linear shadow estimation. But here we evaluate the scalar function
@@
with $Z_{k}=\bigotimes_{l\in[n]}Z^{k_{l}}$ and $Z$ the Pauli-$Z$ operator. In contrast to the scalar function $o$, $\phi$ does not use the dual frame. Assuming the initial state |$\outerproduct{0}{0}$ | can be accurately prepared, the mean of $\phi_{k}$ evaluated on $\mathcal{S}$ is an unbiased estimator for $f_{k}$.
Figure 1: Hardware-efficient circuits for robust shallow shadows. Given an initial state $\varrho$, the circuit implementing the measurement consists of $(D+1)$ random layers. For $D=0$, only a layer of $n$ single-qubit rotations is present. Each extra layer is composed of a sub-layer of entangling gates followed by $n$ single-qubit rotations. At the end, a computational basis measurement is performed.
@@ Results.
The starting observation for our work is the following: Let $\nu$ be an arbitrary random ensemble of quantum circuits whose implementation suffers time-stationary, gate-dependent noise and errors. The noise can be non-Markovian. If the first gate layer of the circuits, that is applied to the state under scrutiny, accurately implements local channel twirls, then $\tilde{S}$ in Eq. (3) still has the structure of Eq. (4). Hence, $\tilde{S}$ can be estimated via the above calibration procedure introduced for local Clifford rotations in Ref. [27]. A practically motivated assumption is that the major source of noise in the implemented circuits are entangling gates and readout noise. If the first layer then consist of accurately implemented Haar random unitary or Clifford single qubit gates, the invariance assumption holds.
@@
If the quantum circuit implementing $g$ has low depth, we can make use of tensor network methods at multiple places of the protocol in order to reduce the runtime and storage complexity of the calibration protocol as well as partially for the estimation protocol. To be concrete, we specify a circuit ensemble that is directly motivated by existing hardware architectures and applications. The circuit architecture we use is based on the so-called hardware-efficient ansatz [36], depicted in Fig. 1. The circuit starts with a layer of single-qubit rotations, i.i.d. sampled from the Haar measure on single-qubit unitaries or Clifford unitaries. This layer is labelled as depth $D=0$. Every subsequent layer is composed of a “sub-layer” of entangling gates followed by another sub-layer of single-qubit rotations sampled from the same measure as before. In principle, for the subsequent single qubit layers also other measures are admissible.
The entangling layer of the hardware-efficient ansatz has a matrix-product operator (MPO) representation of bond dimension $\chi_{C}=2$, see App. C. Thus, $\Pi_{z,g}$ for $g$ of depth $D$ is exactly described by a matrix-product state (MPS) — also known as tensor train (TT) — of bond dimension (TT-rank) $\chi_{C}^{D}$. As a consequence, also the vector of scalar functions $\phi(z,g)=(\phi_{k}(z,g))_{k\in[n]}$ is described by a TT of rank at most $\chi_{C}^{2D}$ and can be efficiently computed and stored for every observed tuple $(z,g)$.
Furthermore, if the noise does not introduce sizeable non-local correlation, also $\tilde{S}$ has an efficient tensor representation and the coefficient vector $f=(f_{k})_{k\in[n]}\in\mathbb{R}^{n}$ is given by a low-rank TT. We illustrate this argument with a specific noise model in App. C. In particular, for single-qubit Pauli noise, the TT-rank of $f$ is at most $\chi=4^{D}$. This justifies to formulate the post-processing of the calibration procedure as the TT recovery problem
$f=\operatorname*{arg\,min}_{\text{{\hbox{f'}} rank-{\hbox{\chi}} TT}}\left\|\frac{1}{|\mathcal{S}|}\sum_{(z,g)\in\mathcal{S}}\phi(z,g)-f^{\prime}\right\|_{\ell_{2}}\,,$ (6)
@@
Standard approaches for tensor recovery such as modified alternating least-square optimization (MALS) [28] can be readily adopted. In particular, for MALS, we can perform the local update step for every core using contractions and pseudo-inverses with storage and time complexity scaling polynomial in the system size, TT ranks, and number of samples. This suggests that for sufficiently local noise and shallow circuits $\tilde{S}$ can be efficiently learned. Below, we take this as a motivation to numerically investigate the performance of the scheme using MALS. Note, however, that showing overall efficiency would additionally require to prove a guarantee for the TT recovery algorithms from $\mathcal{S}$ with polynomial cardinality.
The construction of the dual frame for the estimation protocol requires the inverse of $\tilde{S}$. Since Eq. (4) is an orthogonal decomposition, $\tilde{S}^{-1}=\sum_{k\in\{0,1\}^{n}}(f_{k})^{-1}\Phi_{k}$. Even though $f$ is a low-rank TT, its element-wise inverse $f^{-1}$ might have higher rank. Nonetheless, one can attempt to find a low-rank TT description either from an estimate for $f$ or directly from the data using the TT recovery algorithms.
Figure 2: Numerical experiments under two-qubit incoherent noise model. (top) Worst-case bias for the estimation of Pauli-observables without mitigation for different system sizes, depth and noise levels. Single-qubit layers sampled from the Haar measure. Noise level indicated by $r$, the average gate infidelity per CNOT, and samples |$\absolutevalue{\mathcal{S}}=10^{6}$. We observe a sizeable bias about an order of magnitude larger than the two-qubit infidelity and moderately increasing with the system size. (bottom) Error in fidelity estimation due to finite statistics $|\mathcal{S}|=10^{5}$ for different system sizes, using both $\tilde{S}^{-1}$ and $S^{-1}$. Single-qubit layers sampled from the Haar measure. The error is averaged over $50$ perfectly prepared uniformly-random pure states; errors bars denote the standard error. Noise only affects CNOT gates with $r=10^{-3}$. We observe a reduction of the system-size dependence and, in consequence, smaller estimation errors for shallow circuit compared to (noiseless) local bases measurements ($D=0$). The error in unmitigated fidelity estimation grows linearly with $n$, with a higher slope for increasing depth. In contrast, for mitigates fidelity estimation the slope decreases for increasing depth. (inset) Fidelity estimation error using $\tilde{S}^{-1}$ as a function of $r$ for $n=10$ qubits and $|\mathcal{S}|=10^{5}$. We observe a cross-over point at $r=10^{-2}$; for smaller (larger) infidelities the error decreases (increases) with circuit depth.
@@
Here, we numerically study the performance of robust estimation with shallow circuits using architectures that are motivated by applications. We use two noise models: (i) two-qubit incoherent noise model, where a different unitary noise is applied to a circuit at each execution (see App. A for details); (ii) a composition of a two-qubit Pauli noise with gate-dependent single-qubit Pauli noise, where different single-qubit gates suffer from different Pauli noise channels.

### v2

# Robust ultra-shallow shadows

@@ Abstract
We present a robust shadow estimation protocol for wide classes of low-depth measurement circuits that mitigates noise as long as the effective measurement map including noise is locally unitarily invariant. This is in practice an excellent approximation, encompassing for instance the case of ideal single-qubit Clifford gates composing the first circuit layer of an otherwise arbitrary circuit architecture and even non-Markovian, gate-dependent noise in the rest of the circuit. We argue that for weakly-correlated local noise, the measurement channel has an efficient matrix-product representation, and show how to estimate this directly from experimental data using tensor-network tools, eliminating the need for analytical or numeric calculations. We illustrate the relevance of our method with both numerics and proof-of-principle experiments on an IBM Quantum device. Numerically, we show that unmitigated shallow shadows with noisy circuits become more biased as the depth increases. In contrast, using the same number of samples, robust ultra-shallow shadows become more precise with increasing depth for relevant parameter regimes. The gain in sample efficiency is still limited by the noise per gate, resulting in an optimal circuit depth per noise level. Experimentally, we observe improved precision in two simple fidelity estimation tasks using five-qubit circuits with up to two layers of entangling gates, by about an order of magnitude. Under the practical constraints of current and near-term noisy quantum devices, our method maximally realizes the potential of shadow estimation with global rotations and identifies its fundamental limitations in the presence of noise.
@@ Introduction.
Randomized measurements are a central tool for the estimation of observables on repeatedly prepared quantum states [1, 2, 3]. Applications abound [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17]. Ref. [18] derived guarantees and proved optimality of linear estimation from bases measurements generated by global or local Clifford rotations and coined the term “classical shadows”. Besides optimality, an important contribution of Ref. [18] is the efficiency of estimation from global Clifford rotations. For example, with the same classical shadow from constantly many state copies in the system size $n$, one can estimate many different pure state fidelities. In practice, however, harnessing this efficiency is typically not possible, since global Clifford rotations require the accurate implementation of “long” circuits of ($\order{n^{2}/\log(n)}$ ) many local gates [19]. This obstacle has motivated the study of “shallow shadows”, where shorter circuits are employed to “interpolate” between local and global rotations [20, 21, 22, 23, 24]. There is evidence that for certain observables classical shadows with global rotations can be approximately de-randomized to $\mathcal{O}(n\log(n))$ gates [24]. On noisy intermediate-scale quantum (NISQ) hardware, however, even measurement circuits in this scaling regime are challenging. For the measurement, one typically wants to work with hardware-efficient circuits of very low depth. For clarity, we refer to such construction as ultra-shallow shadows.
@@
The two-design property of uniformly random Clifford rotations allows for straight-forward analytical inversion of the effective measurement map [25], a prerequisite to constructing the classical shadow. However, the works on shallow shadows resort to numerics for the evaluation of the effective measurement map for depth larger than two [23], exploiting the local invariance of the circuit ensembles [20, 21, 22, 24]. However, for measurement circuits involving entangling gates, noise, and (calibration) errors will typically cause sizable biases in the derived estimators on current NISQ hardware. For this reason, using ultra-shallow circuits for precise estimations in practice unavoidably requires a variant of classical shadows that mitigates circuit noise. That such mitigation is possible for classical shadows with single-qubit or global random Clifford rotations if the noise is well-characterized was shown for simple depolarizing noise models [26], local gate-dependent noise models [27], and via randomized benchmarking experiments [28]  [^1] Note that related but differently motivated research lines are mitigating noise in the state preparation using non-linear shadow estimation [40, 41, 9, 42, 43, 44, 45], or existing error-mitigation techniques [46], and the optimization of dual frames [47], e.g.. Going beyond these works, ideally one wishes for a robust shadow scheme that replaces analytical or numerical calculations of the measurement map with a direct empirical estimation of it in the presence of noise. Such a tool was developed in Ref. [30], again for the cases of single-qubit or global random Clifford rotations under the assumptions of gate-independent Markovian noise acting after the rotation. However, a careful extension to the most crucial, experimentally relevant regime of ultra-shallow circuits, where no analytical calculation of the measurement frame is, in general, available, is missing. Despite the large body of literature on classical shadows, a central question is still open: Can the theoretically established improvement in sample efficiency of randomizing measurements with entangling gates be turned into a practical advantage in near-term experiments?
@@
Our work answers this question in the affirmative. We show that achievable advantages are significant but also limited at current noise levels. To this end, we present a robust shadow estimation scheme for ultra-shallow circuits with arbitrary architectures of random local gates, under mild and practically plausible assumptions on the noise. We assume only that the average noisy measurement channel is invariant under local rotations and the accurate implementation of a single initial state. A simple but experimentally crucial observation is that, for this to happen, it is sufficient for instance that the first circuit layer consists of accurately implemented single-qubit random Cliffords, regardless of the subsequent circuit, which can even be subject to time-stationary, gate-dependent, non-Markovian noise. As observed in previous works, single-qubit Pauli invariance guarantees that the measurement map is Pauli diagonal and the circuit shallowness guarantees that the corresponding diagonal elements admit a tensor train (TT) description of low-rank. For us, it is crucial that this structure is maintained if the noise is approximately local. This allows us to develop efficient methods for directly estimating the measurement channel using standard tensor-network tools such as the modified alternating least squares ($\operatorname{MALS}$) [31].
@@
We numerically quantify how unmitigated shadow estimations (of both fidelities and Pauli observables) in the presence of noise become increasingly biased as the circuit depth increases. In contrast, we show that, for experimentally relevant regimes, robust ultra-shallow shadows become more precise as the depth increases up to an optimal depth depending on the strength of the incoherent noise. For instance, for depth $D=2$ and $n=7$, we observe that with our method average errors of fidelity estimation with random target states improve by a factor of two over noise-free local shadows. Without mitigation, errors of shallow shadows are about an order of magnitude larger than with our method. We find that, even for these small depths and system sizes, infidelities of the entangling gates of at most $10^{-3}$ are required to realize improvements in precision at all. Moreover, we apply our scheme to a recent experiment [5] where ideal, four-qubit global Clifford rotations were approximated by noisy circuits of depth five, in an unmitigated fidelity estimation. In numerical simulations, we observe a four-fold reduction of error with our method. Finally, we perform a proof-of-principle experiment on the ibm_nairobi device. We estimate the effective measurement map and find that, indeed, simplified assumptions beyond what our methods exploit are not permissible.
Our framework offers a versatile and practical solution to classical shadows in the most relevant regime for near-term quantum devices, where randomized measurements are implemented by noisy shallow circuits.
@@
We start by reviewing the classical shadow protocol on an ideal noiseless device. Consider a quantum device that can be repeatedly prepared in an unknown $n$-qubit quantum state $\varrho$. Furthermore, we will work with a quantum device that is capable of implementing circuits composed from a unitary gate set $\mathcal{G}$ and of measuring in a fixed (computational) basis, specified by the effects =$\mathcal{B}=\{\Pi_{z}=\outerproduct{z}{z}\}_{z\in[d]}$, with $d=2^{n}$ and $[d]=\{0,1,\dots,d-1\}$.
@@
Every iteration of the protocol is described by the same positive operator-valued measure (POVM) taking values $\Pi_{z,g}=g^{\dagger}\,\Pi_{z}\,g$ with $g$ in the support of $\nu$ and $z\in[d]$. If the support of $\nu$ is sufficiently large, the POVM is informationally complete (IC), and asymptotically any information about $\varrho$ can be reconstructed from the measurement statistic $\mathcal{S}$.
@@
where we have used Dirac notation w.r.t. the Hilbert-Schmidt inner product. Since the POVM is informationally complete (IC), $S$ is invertible and the elements of the canonical dual frame are given by $\bar{\Pi}_{z,g}=S^{-1}(\Pi_{z,g})$, fulfilling $\mathbb{E}_{g\sim\nu}{\big{[}}\sum_{z\in[d]}|\bar{\Pi}_{z,g})(\Pi_{z,g}|{\big{]}}=S^{-1}S=\mathbbm{1}_{d^{2}}$, with $\mathbbm{1}_{d^{2}}$ being the $d^{2}\times d^{2}$ identity operator. In the classical post-processing, we evaluate the scalar function $o(z,g)=(O|\bar{\Pi}_{z,g})$ on each entry of $\mathcal{S}$ and take the empirical mean $\hat{o}=\sum_{(z,g)\in\mathcal{S}}o(z,g)$. For the ideal device, we observe $(z,g)$ with probability density $p(z,g)=(\Pi_{z,g}|\varrho)\,\mathrm{d}\nu(g)$ and, thus, $\hat{o}$ is indeed an unbiased estimator of
@@
In practice, however, experimental imperfections in the implementation of the basis measurement $\mathcal{B}$ and the circuits in the support of $\nu$ induce a bias in the estimator. We denote by $\tilde{\Pi}_{z,g}$ the elements of the POVM that are implemented on the device and define the “one-sided” noisy frame operator as
@@
The same argument above now yields a non-vanishing estimator bias such that $\operatorname{bias}(O,\varrho)=|\mathbb{E}(\hat{o})-(O|\varrho)|=|(O|S^{-1}\tilde{S}-\mathbbm{1}_{d^{2}}|\varrho)|$.
The form of the bias already hints at a conceptually simple solution to mitigate the noise. Assume we have an empirical estimate of the noisy frame operator $\tilde{S}$ and it is still invertible, which happens if $\{\tilde{\Pi}_{z,g}\}_{z,g}$ is still informationally complete. We can make use of this estimate in the construction of a scalar function $\tilde{o}(z,g)=(O|\tilde{S}^{-1}|\Pi_{z,g})$ that we evaluate instead of $o(z,g)$ in the classical post-processing. The resulting mean estimator $\hat{\tilde{o}}$ is again unbiased with $\mathbb{E}(\hat{\tilde{o}})=(O|\tilde{S}^{-1}\,\tilde{S}|\varrho)$.
@@
More concretely, recall that conjugation with the Clifford group (as with the unitary group) acts irreducibly on the space of traceless matrices, e.g.  [2]. Correspondingly, tensor-products of single-qubit Clifford channels decompose into $d$ irreducible representations (irreps) without multiplicities, labeled by $k\in\{0,1\}^{n}$. The projectors onto the irreps are $\Phi_{k}=\bigotimes_{l\in[n]}\Phi_{k_{l}}$ with $\Phi_{0}=d^{-2}\,|\mathbbm{1}_{2})(\mathbbm{1}_{2}|$ and $\Phi_{1}=\mathbbm{1}_{4}-\Phi_{0}$. Assuming single-qubit Clifford-invariance, the noisy frame operator can, thus, be written as
@@
Ref. [30] proposed the following calibration procedure to learn the coefficients $f_{k}$: Perform a randomized bases measurement on the input state |$\outerproduct{0}{0}$ |. The post-processing is analogous to that of linear shadow estimation. But here we evaluate the scalar function
@@
with $Z_{k}=\bigotimes_{l\in[n]}Z^{k_{l}}$ and $Z$ the Pauli-$Z$ operator. In contrast to the scalar function $o$, $\phi$ does not use the dual frame. Assuming the initial state |$\outerproduct{0}{0}$ | can be accurately prepared, the mean of $\phi_{k}$ evaluated on $\mathcal{S}$ is an unbiased estimator for $f_{k}$. From the form of $\tilde{o}(z,g)$ and Eq. (5), we see that the spectrum of $\tilde{S}$ must be characterized to relative precision to guarantee controlled additive precision in estimator $\hat{\tilde{o}}$.
Figure 1: Hardware-efficient circuits for robust shallow shadows. Given an initial state $\varrho$, the circuit implementing the measurement consists of $(D+1)$ random layers. For $D=0$, only a layer of $n$ single-qubit rotations is present. Each extra layer is composed of a sub-layer of entangling gates followed by $n$ single-qubit rotations. At the end, a computational basis measurement is performed. We assume that the implementation of the circuit can suffer from noise that does not break the local invariance and low-rank representation of the measurement map. For instance, approximately local gate-dependent non-Markovian noise (indicated with the lightning symbols) affecting the gates after the first gate-layer and the read-out is admissible.
@@ Results.
The starting observation for our work is the following: Let $\nu$ be an arbitrary random ensemble of quantum circuits whose implementation suffers time-stationary, gate-dependent noise and errors. The noise can be non-Markovian. If the first gate layer of the circuits, that is applied to the state under scrutiny, accurately implements local channel twirls, then $\tilde{S}$ in Eq. (3) still has the structure of Eq. (4) even in the presence of noise. Hence, $\tilde{S}$ can be estimated via the above calibration procedure introduced for local Clifford rotations in Ref. [30]. A practically motivated assumption is that the major sources of noise in the implemented circuits are entangling gates and readout noise. If the first layer then consists of accurately implemented Haar random unitary or Clifford single qubit gates, the invariance assumption holds.
@@
If the quantum circuit implementing $g$ has low depth, we can make use of tensor network methods at multiple places of the protocol to reduce the runtime and storage complexity of the calibration protocol as well as partially for the estimation protocol. To be concrete, we specify a circuit ensemble that is directly motivated by existing hardware architectures and applications. The circuit architecture we use is based on the so-called hardware-efficient ansatz [35], depicted in Fig. 1. The circuit starts with a layer of single-qubit rotations, i.i.d. sampled from the Haar measure on single-qubit unitaries or Clifford unitaries. This layer is labeled as depth $D=0$. Every subsequent layer is composed of a “sub-layer” of entangling gates followed by another sub-layer of single-qubit rotations sampled from the same measure as before. In principle, for the subsequent single qubit layers also other measures are admissible.
The entangling layer of the hardware-efficient ansatz has a matrix-product operator (MPO) representation of bond dimension $\chi_{C}=2$, see App. A. Thus, $\Pi_{z,g}$ for $g$ of depth $D$ is exactly described by a matrix-product state (MPS)—also known as tensor train (TT)—of bond dimension (TT-rank) $\chi_{C}^{D}$. As a consequence, also the vector of scalar functions $\phi(z,g)=(\phi_{k}(z,g))_{k\in[n]}$ is described by a TT of rank at most $\chi_{C}^{2D}$ and can be efficiently computed and stored for every observed tuple $(z,g)$.
Furthermore, if the noise does not introduce sizable non-local correlation, also $\tilde{S}$ has an efficient tensor representation, and the coefficient vector $f=(f_{k})_{k\in[n]}\in\mathbb{R}^{n}$ is given by a low-rank TT. We illustrate this argument with a specific noise model in App. A. In particular, for single-qubit Pauli noise, the TT-rank of $f$ is at most $\chi=4^{D}$. This justifies formulating the post-processing of the calibration procedure as the TT recovery problem
$f=\operatorname*{arg\,min}_{\text{$f^{\prime}$ rank-$\chi$ TT}}\left\|\,\frac{1}{|\mathcal{S}|}\sum_{(z,g)\in\mathcal{S}}\phi(z,g)-f^{\prime}\,\right\|_{\ell_{2}}\,,$ (6)
@@
Standard approaches for tensor recovery such as modified alternating least-square optimization (MALS) [31] can be readily adopted. In particular, for MALS, we can perform the local update step for every core using contractions and pseudo-inverses with storage and time complexity scaling polynomially in the system size, TT ranks, and number of samples. This suggests that for sufficiently local noise and shallow circuits, $\tilde{S}$ can be efficiently learned. Below, we numerically investigate the performance of the scheme using MALS. Note, however, that analytically establishing the overall efficiency would additionally require proof of a guarantee for the TT recovery algorithms from $\mathcal{S}$ with polynomial cardinality.
The construction of the dual frame for the estimation protocol uses the inverse of $\tilde{S}$. Since Eq. (4) is an orthogonal decomposition, $\tilde{S}^{-1}=\sum_{k\in\{0,1\}^{n}}(f_{k})^{-1}\Phi_{k}$. From $f$ in TT format, we can efficiently evaluate entries of its element-wise inverse $f^{-1}$ and, thus, the eigenvalues of $\tilde{S}^{-1}$. For certain observables that themselves have an efficient tensor-network representation, having $f^{-1}$ in low-rank TT format would even allow for efficiently computing the estimators $\hat{\tilde{o}}$. Even though $f$ is a low-rank TT, $f^{-1}$ might have a higher rank. Finding a low-rank MPS representation for the inverse of a low-rank MPS is, in general, an open question. Nonetheless, one can attempt to find a low-rank TT description either from the estimate for $f$ or directly from the data using the TT recovery algorithms.
Figure 2: Numerical experiments under two-qubit incoherent noise model. (top) Worst-case bias for the estimation of Pauli-observables without mitigation for different system sizes, depth, and noise levels. Single-qubit layers sampled from the Haar measure. Noise level indicated by $r$, the average gate infidelity per CNOT, and samples |$\absolutevalue{\mathcal{S}}=10^{6}$. We observe a sizable bias about an order of magnitude larger than the two-qubit infidelity and moderately increasing with the system size. (bottom) Error in fidelity estimation due to finite statistics |$\absolutevalue{\mathcal{S}}=10^{5}$ for different system sizes, using both $\tilde{S}^{-1}$ and $S^{-1}$. Single-qubit layers sampled from the Haar measure. The error is averaged over $50$ perfectly prepared uniformly random pure states; errors bars denote the standard error. Noise only affects CNOT gates with $r=10^{-3}$. We observe a reduction of the system-size dependence and, in consequence, smaller estimation errors for shallow circuits compared to (noiseless) local basis measurements ($D=0$). The error in unmitigated fidelity estimation grows linearly with $n$, with a higher slope for increasing depth. In contrast, for mitigated fidelity estimation the slope decreases with increasing depth. (inset) Fidelity estimation error using $\tilde{S}^{-1}$ as a function of $r$ for $n=10$ qubits and |$\absolutevalue{\mathcal{S}}=10^{5}$. We observe a cross-over point at $r=10^{-2}$; for smaller (larger) infidelities the error decreases (increases) with circuit depth.
@@
Here, we numerically study the performance of robust estimation with shallow circuits using architectures that are motivated by applications. We use two noise models: ($i$) two-qubit incoherent noise model, where a different unitary noise is applied to a circuit at each execution (see App. B for details); ($ii$) a composition of a two-qubit Pauli noise with gate-dependent single-qubit Pauli noise, where different single-qubit gates suffer from different Pauli noise channels.